OpenStudy (anonymous):

Compare and Contrast: Two equations are listed below. Solve each equation and compare the solutions. Choose the statement that is true about both solutions. Equation 1 Equation 2 |5x + 6| = 41 |2x + 13| = 28 Equation 1 has more solutions than equation 2. Equation 1 and Equation 2 have the same number of solutions. Equation 2 has more solutions than Equation 1. The number of solutions cannot be determined.

OpenStudy (anonymous):

FIRST EQUATION: ´´´´´´´´´´´´´´ |5x + 6| = 41 has two valid solutions for -41 and 41 which both yield abs 41 |5x+6| = |41| = 41 and |5x+6| = |-41| = 41 solve these two solutions: a) 5x+6 = 41 5x = 35 x = 7 b) 5x+6 = -41 5x = -47 x = -9 -2/5

OpenStudy (anonymous):

Solution One correct Solution Two -41

OpenStudy (anonymous):

I think they both have two solutions, they would usually have one because you can solve for x. However because abs also accepts the negative counterpart of the solution equally well, we have 2 solutions for every eq

OpenStudy (anonymous):

Thanks